Tatianaw:

George and Carmen went on bicycle trip. They took a bus to their starting points, and then biked the rest. They traveled 350 kilometers in total, and they biked 80 kilometers more than they were bussed. How many kilometers they traveled by bike?

KyledaGreat:

Let's call the distance that George and Carmen traveled by bus "x". We know that they traveled a total of 350 kilometers, so the distance they traveled by bike must be: 350 - x We also know that they biked 80 kilometers more than they were bussed, so: distance biked = distance bussed + 80 We can set up an equation based on this information: 350 - x = x + 80 Simplifying the equation, we get: 270 = 2x + 80 Subtracting 80 from both sides, we get: 190 = 2x Dividing by 2, we get: x = 95 So George and Carmen traveled 95 kilometers by bus, and 350 - 95 = 255 kilometers by bike. Therefore, they traveled 255 kilometers by bike.

Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment
Dolphan:

1 attachment